scheduled payments of $547, $596 and $262 are due in one and a half years, four years and five and a half years respectively. what is the single replacement payment 2 and a half years from nowif interest is 5.1% compounded monthly?
Answer #1

Computation of value of $ 547 on year two-and-half-years which is schedule on year one-and-half years:

FV = PV x (1+r) n

PV = $ 547

r = Periodic rate = 0.051/12 = 0.00425

n = 1-year x 12 periods = 12 periods

Value of $ 547 on two-and-half-years from now

= $ 547x (1+ 0.00425) 12

= $ 547 x (1.00425) 12

= $ 547x 1.05220917603682

= $ 575.558419292141

Computation of Value of $ 596 on year two and -half-year which is schedule on four years:

PV = FV/(1+r) n

FV = $ 596

r = Periodic rate = 0.051/12 = 0.00425

n = 1.5-years x 12 periods = 18 periods

Value of $ 596 on two-and-half-years from now

= $ 596/ (1+0.00425) 18

= $ 596/ (1.00425) 18

= $ 596/ 1.07932721357736

= $ 552.195842468009

Computation of Value of $ 262 on year two and -half-year which is schedule on five-and-half years:

PV = FV/(1+r) n

FV = $ 262

r = Periodic rate = 0.051/12 = 0.00425

n = 3-years x 12 periods = 36 periods

Value of $ 262 on year two and -half-years from now

= $ 262/ (1+0. 00425) 36

= $ 262/ (1.00425) 36

= $ 262/ 1.16494723396867

= $ 224.902890328718

Total value of three cash flows on year two-and-half-years from now

= $ 575.558419292141 + $ 552.195842468009 + $ 224.902890328718

= $ 1,352.657152088868 or $ 1,352.66

The equivalent replacement payment on two and half years from now is $ 1,352.66